You want to obtain a sample to estimate a population proportion.
Based on previous evidence, you believe the population proportion is 58%.
You would like to be 99% confident that your estimate is within 2.5% of the true population proportion.
How large of a sample size is required?
Do not round mid-calculation.
n=

Answer by Boreal(15235)

error=z(2.5768)*sqrt(0.58*0.42/n)
0.025^2=z^2*0.58*0.42/n
n=z^2*0.58*0.42/0.025^2
=2586.01 or 2587